# PointSAM: Pointly-Supervised Segment Anything Model for Remote Sensing Images

### v1

@@ III-C Prototype-based Regularization
General self-training methods are prone to confirmation bias [52]. There are two common solutions to solve this problem. The first approach [23, 24, 25] involves aligning the predicted features extracted by the model from the source data with those extracted from the target data. However, due to the vast amount of data used to train SAM, it is challenging to obtain an accurate source distribution. Additionally, the limited number of batches used in SAM’s fine-tuning can also result in inaccurate prediction distributions. Therefore, this approach is not suitable for our task. The second approach [26, 16] introduces an anchor model to obtain the corresponding logits to constrain the predicted logits. Specifically, this method uses the frozen weights of the source model to predict the results on the target data and constrains the self-training process of the target model with these results. However, since the source model’s predictions might contain significant errors, this approach may not deliver the expected results. Instead of directly constraining logits, we propose instance-level constraints without relying on source data. First, we generate target prototypes using GT points through Offline Prototype Generation. Predicted prototypes are then dynamically obtained via Memory Bank Updating. Finally, Hungarian Matching is used to align the target and predicted prototypes. The following sections will provide a detailed explanation of each step in the proposed approach
@@ III-C1 Offline Prototype Generation
We begin by using the source model to extract embeddings offline for prompts corresponding to each instance in the target dataset. As illustrated in Fig. 3(a), given an image $\mathcal{I}_{img}\in\mathbb{R}^{C\times H\times W}$ from the target dataset, we pass it through the frozen image encoder of SAM to obtain the backbone feature map $\mathcal{F}^{b}\in\mathbb{R}^{C_{b}\times H/s\times W/s}$. To locate the ground truth (GT) prompt $\left(x_{k},y_{k}\right)$ within the original image, we map it to the feature map coordinates ($x_{k}^{\prime}={x_{k}}/s$, $y_{k}^{\prime}=y_{k}/s$) and extract the corresponding embedding $f_{k}^{t}\in\mathbb{R}^{C_{b}}$:
@@
In this way, we can obtain a large number of feature points $\{f_{k}^{t}\}_{k=1\cdots K}$ from the source model corresponding to GT points in the target data. Next, we cluster these feature points. Since SAM is a class-agnostic segmenter, the feature points lack class labels, and the number of clusters is unknown. Consequently, using Kmeans for clustering directly is suboptimal. To address this, we employ FINCH [27], a clustering algorithm that does not require prior knowledge of the number of clusters. We then compute the mean feature of each cluster to represent the target prototype. Let $C^{t}_{i}$ denote the $i$-th cluster:
@@
Figure 4: The process of negative prompt calibration. The positive and negative prompts are represented by red points () and green points (), respectively. Different prompts input into SAM generates different initial masks. To refine these masks, an IoU matrix is calculated for each instance pair. Matrix values greater than 0 indicate that the two objects can act as negative constraints for each other. By using the positive prompt of one object as the new negative prompt for another and inputting it into SAM again, a refined mask is generated.
@@ III-C2 Memory Bank Updating
Note that $\mathcal{B}$ is initialized as an empty set $\mathcal{B}=\emptyset$ at the beginning of training. We fill the queue with features using the teacher model without dequeue until $\mathcal{B}$ reaches the predefined length. The update mechanism follows a first-in, first-out (FIFO) approach to dynamically update the predicted feature information and prevent the features from becoming outdated.
Similar to the generation of target prototypes, we also utilize the FINCH [27] algorithm to cluster the features in the memory bank =$\mathcal{B}=\left\{f_{1}^{p},f_{2}^{p},\cdots,f_{K}^{p}\right\}$. Suppose $C^{p}_{j}$ represents the $j$-th cluster in $\mathcal{B}$, the predicted prototypes $\mathcal{P}^{p}_{j}$ can be defined as:
@@ III-C3 Hungarian Matching
Since the target prototypes $\{\mathcal{P}^{t}_{i}\}_{i=1\cdots I}$ and predicted prototypes $\{\mathcal{P}^{p}_{j}\}_{j=1\cdots J}$ cannot be definitively matched one-to-one in order, it is not feasible to use a simple metric function to enforce their consistency. Inspired by the instance matching in DETR [64], we use Hungarian Matching here for computing feature similarity. Then we will describe the specific process.
Define a distance matrix $\mathbf{D}\in\mathbb{R}^{K_{t}\times K_{p}}$, where each element $\mathbf{D}_{ij}$ represents the distance between the $i$-th target prototype $\mathcal{P}^{t}_{i}$ and the $j$-th predicted prototype $\mathcal{P}^{p}_{j}$. Typically, Euclidean distance or cosine similarity is used to compute this distance. Here, we use the latter one as an example to illustrate:
@@
Then the Hungarian algorithm [65] is employed to find the optimal matching that minimizes the total distance. This algorithm solves the bipartite matching problem, finding the permutation $\pi$ that minimizes the total matching loss:
@@
where $\Pi$ represents the set of all possible matchings, and $\pi(i)$ is the index of the predicted prototype matched to the $i$-th target prototype. The final match loss can be expressed as the sum of distances for all matches:
@@ III-D Negative Prompt Calibration
For training SAM, point prompts typically include both positive and negative prompts, which require human annotation. Positive prompts are usually selected from any point in the instance, while negative prompts are more ambiguous. This is because the background often covers a large area, so any point outside the mask can be chosen as a negative prompt. Remote sensing images are characterized by densely packed objects and their high similarity to the background. With point supervision and no boundary constraints, the self-training process may result in a single predicted mask containing multiple foreground objects or extensive background. As shown in Fig. 2, introducing negative prompts can effectively separate objects from confusing regions. Inspired by this, we developed a Negative Prompt Calibration (NPC) method to dynamically adjust negative prompts during training.
Fig. 8 illustrates the complete process of NPC. Given a set of initial prompt points $\mathcal{P}_{\text{init}}$, which includes $K$ positive prompts $\mathcal{P}^{\text{pos}}_{\text{init}}$ = $\{\mathbf{p}^{\text{pos}}_{k}\}_{k=1\cdots K}$ and $K$ negative prompts $\mathcal{P}^{\text{neg}}_{\text{init}}$ = $\{\mathbf{p}^{\text{neg}}_{k}\}_{k=1\cdots K}$. For each $i$-th instance, the corresponding initial mask $\mathcal{M}^{\text{init}}_{i}$ is obtained by inputting $\mathcal{P}_{\text{init}}$ and encoder feature into the mask encoder $\Phi_{\text{mask}}$ and prompt encoder $\Phi_{\text{prompt}}$:
@@
Here, we ignore the feature from the image encoder.
For images containing multiple objects, $M_{\text{init}}$ will also contain multiple masks. We first compute the Intersection over Union (IoU) between each pair of masks and construct an IoU matrix $\mathbf{O}$, where each element $\mathbf{O}_{ij}$ represents the IoU between the $i$-th mask $\mathcal{M}_{i}$ and the $j$-th mask $\mathcal{M}_{j}$. To exclude self-correlation, we set the diagonal elements to 0:
@@
For masks that intersect with a given instance mask, we identify corresponding positive prompts as candidate negative prompts. Specifically, for a given mask $\mathcal{M}_{i}$, the set of candidate negative prompts $\mathcal{P}_{\text{neg},i}$ is derived from the positive prompts of masks that intersect with $\mathcal{M}_{i}$:
@@
After obtaining the new negative prompts, we input them along with the initial positive prompts into SAM’s mask prompt to obtain the final refined masks $\mathcal{M}^{r}$:
@@
In this way, the refined mask $\mathcal{M}^{r}$ can be used as a pseudo-label to supervise the mask $\mathcal{M}^{s}$ predicted by the student. We will introduce the loss function for supervision in the next section.
@@ III-E Total Loss
In the original SAM model, there are three loss functions: IoU loss $\mathcal{L}_{\text{IoU}}$, Dice loss $\mathcal{L}_{\text{dice}}$, and Focal loss $\mathcal{L}_{\text{focal}}$. These losses are applied to the GT masks and predicted masks. However, in our case, since there are no GT masks, we use these three losses to supervise the student model’s predictions $\mathcal{M}^{s}$ with the refined masks $\mathcal{M}^{r}$ predicted by the teacher. Additionally, we include the matching loss $\mathcal{L}_{\text{match}}$ to constrain the target and predicted prototypes. The total loss $\mathcal{L}_{\text{total}}$ is defined as:
@@ IV-B Experiment Details
Competing Methods: We evaluate multiple source-free domain adaptation approaches and the latest weakly supervised interactive segmentation methods. Specifically, directly testing the pre-trained model (Direct) with fixed prompt inputs is susceptible to distribution shifts and may not perform well on target datasets with significant shifts. TENT [69] is a basic test-time adaptation method that adapts to the target domain by optimizing an entropy loss. SHOT [22] employs pseudo labels and applies a uniform distribution assumption for source-free domain adaptation. Self-Training [19] was initially developed for semi-supervised learning. We simply adopt a vanilla teacher-student structure without any tricks. Tribe [26] proposed a strong baseline for test-time adaptation under continual and class-imbalanced domain shifts. We adapt it for domain adaptive segmentation by replacing the training losses. DePT [70] inserts visual prompts into a visual Transformer and adjusts these source-initialized prompts solely during the adaptation process without accessing the source data. WeSAM incorporates anchor loss and prompt-based contrastive loss into self-training. Finally, we evaluate our own pointly-supervised interactive segmentation method, referred to as Ours.
TABLE I: Comparison result of different methods on NWPU VHR-10 test set. Colored results represent the best and second-best.
@@
Implementation Details We finetune the LoRA module of the image encoder by Adam optimizer for all experiments. We set the batch size to 1 using an RTX3090 GPU and the learning rate to 0.0005 with a weight decay of 0.0001. We set the low rank of the LoRA module for the image encoder to 4. And, the coefficients $\lambda_{\text{focal}}$ and $\lambda_{\text{match}}$ in Eq. 14 are set to 20 and 0.1, respectively. We apply strong and weak data augmentations for self-training and choices for augmentation follow [16].
TABLE II: Comparison result of different methods on WHU Building test set. Colored results represent the best and second-best.
@@
TABLE III: Comparison result of different methods on HRSID-inshore test set. Colored results represent the best and second-best.
@@
TABLE IV: Ablation studies of the proposed PointSAM on the HRSID-inshore dataset. ST, NPC, and PBR refer to self-training, negative prompt calibration, and prototype-based regularization, respectively.
@@ IV-C Quantitative Evaluations
We conducted quantitative evaluations across three datasets—NWPU VHR-10, WHU, and HRSID-inshore. All comparison methods were reproduced on both SAM [5] and SAM2 [6], and we compared the IoU and F1 scores for different numbers of points, ranging from 1 to 3.
@@ IV-D3 Alternative IoU threshold
We compared different clustering and feature alignment methods, as shown in Fig. 6. KLD constrains the feature mean and variance of the source and target models on the target data using Kullback-Leibler divergence. Kmeans refers to the use of the Kmeans algorithm for feature clustering in PBR, while FINCH is the clustering method used in our work. The results show that FINCH outperforms other methods across various point settings.KLD performs poorly due to insufficient data, leading to inaccurate variance estimation. Kmeans performs slightly worse than FINCH because it requires manually set, fixed clustering centers that are not adaptive to the feature distribution. More importantly, it is more than three times slower than FINCH.
@@
Figure 8: Visualization of the results of negative prompt calibration during training. The positive and negative prompts are represented by red points () and green points (), respectively. The refined mask is generated from the initial mask after applying negative prompt calibration. This calibration effectively guides the negative prompts to more precise regions, enabling the generation of accurate masks.
IV-D5 How about more points?
We validated the results of different methods under an increased number of point prompts. As shown in Fig. 7, simply adding more points does not consistently lead to better performance. This is because increasing the number of points also raises the likelihood of including low-quality points. Such noise can negatively affect the segmentation results of other points. For the Supervised method, the results remained relatively unchanged due to the presence of full-mask constraints. Direct test achieved its best results with two points; however, as the number of points increased, the F1 score gradually decreased. Similarly, Self-training showed a decline in results due to the generation of noisy pseudo-labels. In contrast, our proposed PointSAM maintained stable results, approaching the performance of Supervised. This is because negative prompt calibration effectively corrected the prompts and reduced the impact of inaccurate masks caused by too many points.
Figure 9: Comparison of visualization results for different methods. Rows 1-3 show the results on NWPU VHR-10 with 1 to 3 prompts; similarly, rows 4-6 display the results on WHU, and rows 7-9 show the results on HRSID-inshore. White arrows or zoom-ins in images are used to highlight.

### v2

@@ III-C Prototype-based Regularization
General self-training methods are prone to confirmation bias [55]. There are two common solutions to solve this problem. The first approach [23, 24, 25] involves aligning the predicted features extracted by the model from the source data with those extracted from the target data. However, due to the vast amount of data used to train SAM, it is challenging to obtain an accurate source distribution. Additionally, the limited number of batches used in SAM’s fine-tuning can also result in inaccurate prediction distributions. Therefore, this approach is not suitable for our task. The second approach [26, 16] introduces an anchor model to obtain the corresponding logits to constrain the predicted logits. Specifically, this method uses the frozen weights of the source model (original SAM model) to predict the results on the target data and constrains the self-training process of the target model with these results. However, since the source model’s predictions might contain significant errors, this approach may not be optimal. Instead of directly constraining predicted logits, we propose instance-level constraints without relying on source data. First, we generate target prototypes using GT points through Offline Prototype Generation. Predicted prototypes are then dynamically obtained via Memory Bank Updating. Finally, Hungarian Matching is used to align the target and predicted prototypes.
@@ III-C1 Offline Prototype Generation
We begin by using the source model to extract embeddings offline for prompts corresponding to each instance in the target dataset. As illustrated in Fig. 3(a), given an image $\mathcal{I}_{img}\in\mathbb{R}^{C\times H\times W}$ from the target dataset, we pass it through the frozen image encoder of SAM to obtain the feature map $\mathcal{F}^{b}\in\mathbb{R}^{C_{b}\times H/s\times W/s}$. Given a ground truth (GT) prompt $\left(x_{k},y_{k}\right)$ within the original image, we map it to the feature map coordinates ($x_{k}^{\prime}={x_{k}}/s$, $y_{k}^{\prime}=y_{k}/s$) and extract the corresponding embedding $f_{k}^{t}\in\mathbb{R}^{C_{b}}$ from $\mathcal{F}^{b}$:
@@
In this way, we can obtain sufficient feature points $\{f_{k}^{t}\}_{k=1\cdots K}$ from the source model corresponding to GT points in the target data. Next, we cluster these feature points. Since SAM is a class-agnostic segmenter, the feature points lack class labels, and the number of clusters is unknown. Consequently, directly applying KMeans for clustering is suboptimal. To overcome this limitation, we leverage the FINCH algorithm [27], which does not require prior knowledge of the number of clusters. Subsequently, the mean feature of each cluster is computed to represent the target prototype. Let $C^{t}_{i}$ denote the $i$-th cluster:
@@
Figure 4: The process of negative prompt calibration. The positive and negative prompts are represented by red points () and green points (), respectively. Different prompts input into SAM generates different initial masks. To refine these masks, an IoU matrix is calculated for each instance pair. Matrix values greater than 0 indicate that the two objects can act as negative constraints for each other. By using the positive prompt of one object as the new negative prompt for another and inputting it into SAM again, a refined mask is generated. It is worth noting that Ground Truth here refers to the mask specified by the prompt for a specific instance, not the mask for all instances.
@@ III-C2 Memory Bank Updating
Here, $\mathcal{B}$ is initialized as an empty set $\mathcal{B}=\emptyset$ at the start of training. The memory bank is populated with features from the teacher model without dequeuing until $\mathcal{B}$ reaches its predefined maximum length.
The update process follows a first-in, first-out (FIFO) strategy to dynamically maintain the feature information and prevent stale features from remaining in the memory bank.
Similar to the process of generating target prototypes, we employ the FINCH [27] algorithm to cluster the features in the memory bank =$\mathcal{B}=\left\{f_{1}^{p},f_{2}^{p},\cdots,f_{K}^{p}\right\}$. Let $C^{p}_{j}$ denote the $j$-th cluster in $\mathcal{B}$, the predicted prototypes $\mathcal{P}^{p}_{j}$ are defined as:
@@ III-C3 Hungarian Matching
Since the target prototypes $\{\mathcal{P}^{t}_{i}\}_{i=1\cdots I}$ and predicted prototypes $\{\mathcal{P}^{p}_{j}\}_{j=1\cdots J}$ cannot be directly matched one-to-one in order, a simple metric function is insufficient to enforce their consistency. Inspired by the instance matching strategy in DETR [72], we adopt the Hungarian Matching algorithm to compute feature similarity.
We first define a distance matrix $\mathbf{D}\in\mathbb{R}^{K_{t}\times K_{p}}$, where each element $\mathbf{D}_{ij}$ represents the distance between the $i$-th target prototype $\mathcal{P}^{t}_{i}$ and the $j$-th predicted prototype $\mathcal{P}^{p}_{j}$. In this work, cosine similarity is used to measure the distance:
@@
The Hungarian algorithm [73] is then applied to solve the bipartite matching problem, finding the optimal permutation $\pi$ that minimizes the total distance:
@@
where $\Pi$ denotes the set of all possible matchings, and $\pi(i)$ indicates the index of the predicted prototype matched to the $i$-th target prototype.
The final matching loss is computed as the total distance for all matched pairs:
@@ III-D Negative Prompt Calibration
In SAM training, point prompts include both positive and negative prompts, which require human annotation. Positive prompts are sampled from any point within an instance, while negative prompts are more ambiguous due to the extensive background. Typically, any point outside the mask can serve as a negative prompt. However, remote sensing images present unique challenges with densely packed objects and high similarity to the background. Without boundary constraints, point supervision in self-training may lead to a single predicted mask covering multiple foreground objects or large background regions. As shown in Fig. 2, introducing negative prompts effectively separates objects from ambiguous regions during inference. Inspired by this, we propose a Negative Prompt Calibration (NPC) method that dynamically adjusts negative prompts during training.
Fig. 8 illustrates the full NPC process. Given an initial set of prompt points $\mathcal{P}_{\text{init}}$, which consists of $K$ positive prompts $\mathcal{P}^{\text{pos}}_{\text{init}}$ = $\{\mathbf{p}^{\text{pos}}_{k}\}_{k=1\cdots K}$ and $K$ negative prompts $\mathcal{P}^{\text{neg}}_{\text{init}}$ = $\{\mathbf{p}^{\text{neg}}_{k}\}_{k=1\cdots K}$, the initial mask $\mathcal{M}^{\text{init}}_{i}$ for each instance is generated by feeding $\mathcal{P}_{\text{init}}$ and the encoder features into the mask encoder $\Phi_{\text{mask}}$ and the prompt encoder $\Phi_{\text{prompt}}$:
@@
Here, we omit the image encoder features for simplicity.
For images containing multiple objects, $\mathcal{M}_{\text{init}}$ will also contain multiple masks. We first compute the Intersection over Union (IoU) between each pair of masks and construct an IoU matrix $\mathbf{O}$, where each element $\mathbf{O}_{ij}$ represents the IoU between the $i$-th mask $\mathcal{M}_{i}$ and the $j$-th mask $\mathcal{M}_{j}$. To exclude self-correlation, we set the diagonal elements to 0:
@@
For masks that intersect with a given instance mask, we identify corresponding positive prompts as candidate negative prompts. Specifically, for a given mask $\mathcal{M}_{i}$, the set of candidate negative prompts $\hat{\mathcal{P}}^{\text{neg}}$ is derived from the positive prompts of masks that intersect with $\mathcal{M}_{i}$:
@@
After obtaining the new negative prompts $\tilde{\mathcal{P}}^{\text{neg}}$, we input them along with the initial positive prompts $\mathcal{P}^{\text{pos}}$ into SAM’s mask prompt to obtain the final refined masks $\mathcal{M}^{r}$:
@@
In this way, the refined mask $\mathcal{M}^{r}$ can be used as a pseudo-label to supervise the mask $\mathcal{M}^{s}$ predicted by the student.
@@ III-E Total Loss
In the original SAM model, three loss functions are used: IoU loss $\mathcal{L}_{\text{IoU}}$, Dice loss $\mathcal{L}_{\text{dice}}$, and Focal loss $\mathcal{L}_{\text{focal}}$. These losses are computed between the ground truth (GT) and predicted masks. In our case, since GT masks are unavailable, these losses are used to supervise the student model’s predictions $\mathcal{M}^{s}$ with the refined masks $\mathcal{M}^{r}$ predicted by the teacher. Additionally, we include the matching loss $\mathcal{L}_{\text{match}}$ to enforce alignment between the target and predicted prototypes. The total loss $\mathcal{L}_{\text{total}}$ is defined as:
@@ IV-B Experiment Details
Competing Methods: We evaluate multiple source-free domain adaptation approaches and the latest weakly supervised interactive segmentation methods. Specifically, directly testing the pre-trained model (Direct) with fixed prompt inputs. TENT [77] is a basic test-time adaptation method that adapts to the target domain by optimizing an entropy loss. SHOT [22] employs pseudo labels and applies a uniform distribution assumption for source-free domain adaptation. Self-Training [19] adopt a vanilla teacher-student structure without any tricks. Tribe [26] leverages anchor loss to constrain self-training. DePT [78] inserts visual prompts into a visual Transformer and adjusts these source-initialized prompts solely during the adaptation process without accessing the source data. WeSAM incorporates anchor loss and prompt-based contrastive loss into self-training.
TABLE I: Comparison of different methods on NWPU VHR-10 test set. Best results are bolded, and second-best results are underlined.
@@
Implementation Details We fine-tune the LoRA module of the image encoder using the Adam optimizer across all experiments. Training is performed on an RTX 3090 GPU with a batch size of 1, a learning rate of 0.0005, and a weight decay of 0.0001. The low-rank dimension of the LoRA module is set to 4. The coefficients $\lambda_{\text{focal}}$ and $\lambda_{\text{match}}$ in Eq. 14 are set to 20 and 0.1, respectively. For self-training, we apply both strong and weak data augmentations, following the augmentation strategies described in [16]. Due to the presence of too many instances in some remote sensing images, to save GPU memory, we set the maximum number of training samples per image to 50.
TABLE II: Comparison of different methods on the WHU test set. Best results are bolded, and second-best results are underlined.
@@
TABLE III: Comparison of different methods on the HRSID-inshore test set. Best results are bolded, and second-best results are underlined.
@@
TABLE IV: Ablation studies of the proposed PointSAM on the HRSID-inshore dataset. ST, PBR, and NPC refer to self-training, prototype-based regularization, and negative prompt calibration, respectively.
@@ IV-C Quantitative Evaluations
We conducted quantitative evaluations on three datasets: NWPU VHR-10, WHU, and HRSID-inshore. All comparison methods were reproduced using both SAM [5] and SAM2 [6]. We compared the IoU and F1 scores across different numbers of points, ranging from 1 to 3.
@@ IV-D3 Alternative IoU threshold
We compared different clustering and feature alignment methods, and the results are shown in Fig. 6. KLD constrains the feature mean and variance of the source and target models on the target data using the Kullback-Leibler divergence. Kmeans refers to using the Kmeans algorithm for feature clustering in PBR. Anchor denotes keeping NPC unchanged and replacing PBR with the anchor model from WeSAM[16]. FINCH is the clustering method adopted in this study. The results demonstrate that FINCH outperforms other methods across various point settings. Due to insufficient data, KLD leads to inaccurate variance estimation and performs poorly. Kmeans performs slightly worse than FINCH because it requires manually setting fixed clustering centers, which are not adaptive to the feature distribution. Moreover, its computational speed is over three times slower than FINCH. The performance of the anchor model is inconsistent across the three different prompt quantities, as it is susceptible to inaccurate logit predictions.
@@
Figure 8: Visualization of negative prompt calibration results during training. Positive prompts are marked with red points (), while negative prompts are marked with green points (). The refined mask is obtained by applying negative prompt calibration to the initial mask. This calibration effectively guides the negative prompts to more precise regions, resulting in improved mask predictions.
IV-D5 Alternative fine-tuning methods
We compared the performance impact of different fine-tuning methods, such as LoRA, Adapter, and LayerNorm. LoRA enables efficient fine-tuning by introducing low-rank matrices into the original model, adding only minimal parameters [36]. The Adapter method inserts lightweight modules into the model layers, allowing task-specific learning without modifying the original parameters; in our experiments, we selected the adapter from this work [79]. LayerNorm fine-tunes only the LayerNorm parameters of the original SAM [80]. As shown in Table VI, all fine-tuning methods perform better than direct testing, except for the Adapter in the 1-point case. LoRA outperforms all other fine-tuning methods across all metrics. The Adapter and LayerNorm methods are relatively limited as they are overly focused on local adjustments, which prevents them from achieving global adaptability in complex remote sensing images.
TABLE VI: The impact of different fine-tuning methods on the HRSID-inshore dataset.
Fine-tuning Method 1-Point 2-Point 3-Point
IoU F1 IoU F1 IoU F1
Direct Test 46.56 57.46 37.80 48.34 28.32 37.57
LoRA [36] 56.06 68.38 57.79 70.50 59.37 72.43
Adapter [79] 40.06 53.15 47.83 61.30 47.09 61.27
LayerNorm [80] 49.26 61.39 53.19 66.33 38.01 50.71
IV-D6 How about more points?
We validated the results of different methods under an increased number of point prompts. As shown in Fig. 7, simply adding more points does not consistently lead to better performance. This is because increasing the number of points also raises the likelihood of including low-quality points. Such noise can negatively affect the segmentation results of other points. For the Supervised method, the results remained relatively unchanged due to the presence of full-mask constraints. Direct test achieved its best results with two points; however, as the number of points increased, the F1 score gradually decreased. Similarly, Self-training showed a decline in results due to the generation of noisy pseudo-labels. In contrast, our proposed PointSAM maintained stable results, approaching the performance of Supervised one. This is because negative prompt calibration effectively corrected the prompts and reduced the impact of inaccurate masks caused by too many points.
Figure 9: Comparative results of different methods on the NWPU VHR-10 dataset. Rows 1-3 present the results under 1, 2, and 3 prompts, respectively. Zoomed-in regions in the images are used to highlight the detail.
Figure 10: Comparative results of different methods on the WHU dataset. Rows 1-3 present the results under 1, 2, and 3 prompts, respectively. White arrows in the images are used to highlight the detail.
Figure 11: Comparative results of different methods on the HRSID-inshore dataset. Rows 1-3 present the results under 1, 2, and 3 prompts, respectively.